OpenStudy (goformit100):

solve :- sin D+ sin 2D + sin 3D ........till n terms

OpenStudy (gorv):

@goformit100 ....we can use fourier transform ???

OpenStudy (goformit100):

How ?

OpenStudy (anonymous):

Let S=sin D+sin 2D+ sin 3D+... nterms C=cosD+cos2D+cos 3D+...n ~terms \[C+\iota S=e ^{\iota D}+e ^{2 \iota D}+e ^{3\iota D}+... n~terms.\] \[=\frac{ e ^{\iota D}\left( e ^{\iota n D}-1 \right) }{e ^{\iota D}-1 }\] write \[e ^{\iota D}=CosD+\iota Sin D\] and \[e ^{\iota n D}=Cos~ n D+\iota~ Sin~ n D\] then rationalise and in the end equate imaginary parts.